# AutoGuide: Automated Generation and Selection of State-Aware Guidelines for Large Language Model Agents

### v1

@@ 3.4 Applying State-Aware Guidelines at Test Time
After extracting a set of state-aware guidelines $\mathcal{G}$ from offline experiences, our method employs these guidelines to enhance the decision-making of an LLM agent during testing. At each timestep, AutoGuide summarizes the current test trajectory $\bm{\tau}$ into a state using our state summarization module $\mathcal{M}_{\text{state}}$. Then, we employ the LLM designed for identifying a matching state in Section 3.3 to find a match and retrieves its corresponding guidelines. If there are more than $k$ guidelines related to the current state, we apply the guideline selection module to choose top-$k$ guidelines:
$\begin{split}\mathcal{M}_{\text{select}}(\texttt{state},\mathcal{G}[\texttt{state}],\bm{\tau}),\end{split}$ (4)
@@
where Section B.4 details the prompt template for this selection procedure. Subsequently, AutoGuide incorporates both the state and guidelines into the agent’s prompt. Finally, the agent selects an action by considering the provided state and guidelines (see Figure 1 for an example). This process iterates until the end of the test trajectory (see Algorithm 2 for pseudocode).
@@
Key benefits of AutoGuide. First, the extraction of state-aware guidelines in AutoGuide offers the inherent benefit of providing relevant guidelines for the state of interest. This capability is important since neglecting the specific context in which a guideline applies can confuse the agent’s decision-making process. The second key benefit is the generation of concise natural language guidelines, which can be seamlessly incorporated into any prompt-based LLM agent. Lastly, AutoGuide generates guidelines at the individual state level rather than at the trajectory level. Given that a single incorrect action can lead to a complete failure, it is essential to provide detailed assistance in each action selection process. With these advantages, we demonstrate in the next section that our approach significantly enhances the performance of LLM agents.
@@ 4 Evaluation
We demonstrate the efficacy of AutoGuide by conducting experiments on a diverse suite of sequential decision-making benchmark domains. We also perform important analyses about AutoGuide. Additional experimental details and hyperparameters can be found in Appendix A.
@@ 4.1 Evaluation Setup
4.1.1 Decision-Making Benchmark domains
@@
We consider the following interactive text-based sequential decision-making benchmarks to study different aspects of AutoGuide (see Figure 3):
@@
  • ALFWorld (Shridhar et al., 2021): In this embodied benchmark, an LLM agent interacts with an environment to carry out household tasks, such as placing a pan on the dining table. Observations and actions are expressed in natural language statements, and the agent must navigate through the space and manipulate objects to successfully complete the tasks.
@@
  • WebShop (Yao et al., 2022a): This interactive web environment simulates the task of online shopping on an e-commerce website. The agent’s goal is to understand a text instruction and buy a product that meets specified criteria. This involves querying the website’s search engine, understanding the descriptions and details of each item, and selecting necessary options.
@@
  • WebArena (Zhou et al., 2023): This web-based benchmark introduces realistic environments by replicating the functionality and data found in popular web domains (e.g., Gitlab, Reddit, Wikipedia). Compared to WebShop, WebArena presents more challenges and difficulties for an LLM agent due to its large observation and action space, along with tasks that involve longer planning horizons. We focus on the Reddit domain for the WebArena experiments.
@@
Table 1: Test reward and success rate on WebShop. Agent model for all methods is basedon GPT-3.5-turbo. Reflexion is done by GPT-4-turbo for at most 3 trials.
Table 2: Test result on ALFWorld.
@@
Algorithm WebArena SR
ReAct $8.0$%
ExpeL $21.8$%
AutoGuide $43.7$%
Table 1: Test reward and success rate on WebShop. Agent model for all methods is basedon GPT-3.5-turbo. Reflexion is done by GPT-4-turbo for at most 3 trials.
Table 2: Test result on ALFWorld.
Table 3: Test result on WebArena.
4.1.2 Baselines
We compare AutoGuide against the following baseline approaches to fully show the effect of state-aware guidelines (refer to Appendix A for more details):
  • ReAct (Yao et al., 2022b): This LLM-based planning method integrates reasoning and acting to address sequential decision-making tasks. However, it does not leverage offline experiences and thus suffers from the limited understanding of pre-trained LLMs in downstream domains.
  • ExpeL (Zhao et al., 2023): This method also extracts natural language knowledge from offline data. However, it fails to consider the applicability of guidelines and does not generate state-aware guidelines. Instead, it provides all guidelines to an LLM agent without filtering out irrelevant ones based on the current state. ExpeL has two contributions, the guideline generation and in-context example selection module. Because the latter is orthogonal to our analysis and can be seamlessly combined with our method, we consider ExpeL with guidelines in our experiments.
  • Reflexion (Shinn et al., 2023): This approach converts environmental feedback into text statements to assist an LLM agent (e.g., ReAct) in the next trial. The baseline generates valuable feedback about solving a specific test task. We demonstrate how state-aware guidelines derived by AutoGuide can be combined with the feedback.
4.1.3 Implementation
We collect offline experiences either by running ReAct and Reflexion, or incorporating human demonstrations. We use ReAct with GPT-3.5-turbo as our base LLM agent for WebShop and ALFWorld and GPT-4-turbo for WebArena. For each benchmark, we apply the same GPT version for action generation, state summarization, and guideline selection. We extract state-aware guidelines from the training set with GPT-4-turbo and evaluate their effectiveness by applying them to the test set with non-overlapping tasks. We refer to Appendix A for more details.
Figure 4: Trajectories of ReAct, ExpeL, and AutoGuide from the same test task. ReAct (Left) chose the wrong item, consequently failing the task in the end. ExpeL (middle) was confused by guidelines that were irrelevant to its state, leading to incorrect reasoning and actions. AutoGuide (right) can select useful guidelines relevant to the agent’s state, enabling the agent to efficiently accomplish the task.
4.2 Main Results
Q1. How effective is AutoGuide compared to the competitive baseline methods without state-aware guidelines?
To answer this question, we run ReAct, ExpeL, and AutoGuide on WebShop and ALFWorld benchmarks. The performance on the test datasets is presented in Tables 3 and 3. There are three notable observations:
  1. 1.
    Effectiveness of state-aware guidelines. Our approach surpasses baseline performance in both benchmarks, achieving the highest test rewards and success rates in Tables 3 and 3. These results highlight the effectiveness of employing state-aware guidelines in language-based decision-making domains. To further examine the action selection differences among ReAct, ExpeL, and our method, we present their trajectories in Figure 4. We observed that ReAct makes common mistakes such as trying to take soapbar that is not visible, or taking a soapbottle instead of soapbar due to their similar names. Both ExpeL and AutoGuide improve on this by extracting guidelines from similar mistakes in the offline experience. However, ExpeL often erroneously applies incorrect guidelines due to the availability of all guidelines at each timestep. In Figure 4, ExpeL mistakenly attends to the second guideline “ensure to specify the item’s number and location…”, leading to wrong reasoning and action. AutoGuide presents relevant guidelines at necessary moments, enabling accurate task completion by avoiding the mistakes seen in ExpeL and ReAct.
  2. 2.
    Limitation of in-context learning. In Table 3, while increasing the number of in-context examples (i.e., #-shot) for ReAct gradually improves performance, there is a plateau at a certain number of shots. Additionally, ReAct with more than 6 shots often exceeds the token limit of GPT-3.5-turbo. These results indicate that directly inputting raw trajectories into ReAct for in-context learning is not an effective way to fully leverage offline data. In contrast, AutoGuide extracts knowledge from entire training trajectories by summarizing them into concise state-aware guidelines, making them easy to integrate with prompt-based agents.
  3. 3.
    Importance of providing pertinent knowledge. ExpeL approach helps ReAct by extracting knowledge from offline experiences, but its impact is not as significant as AutoGuide. Recall that for ExpeL, the guidelines are neither generated for specific states at training time, nor selected to only provide state-aware guidelines at test time. As a result, irrelevant guidelines can be introduced to an agent, potentially causing confusion for the agent. Consequently, the result highlights the significance of providing relevant guidelines conditioned on states for LLM agents.
Q2. Does AutoGuide scale to more challenging and complex environments?
In this question, we highlight the scalability of AutoGuide in addressing challenging and complex domains. Specifically, we conduct experiments on WebArena-Reddit, which features more diverse tasks on realistic and complex websites requiring longer action sequences. It has the larger observation space and more complex action space (e.g., scrolling). Table 3 presents the results, where AutoGuide achieves the highest success rate with a significant margin when compared to ReAct and ExpeL. We observe that ReAct scores low task success rate (8.0%) in WebArena due to the complex observation and action spaces and longer task horizon. In ExpeL, the issue of presenting all guidelines to an agent is exacerbated in the WebArena compared to simpler environments like ALFWorld and WebShop. WebArena’s wide variety of tasks across different domains requires a larger number of guidelines to cover the knowledge needed for all tasks and domains. This results in either an overload of irrelevant guidelines that could mislead the agent or a lack of crucial information when the number of guidelines is limited, as suggested in ExpeL (Zhao et al., 2023). In contrast, AutoGuide achieves a more significant performance enhancement (43.7%) compared to ExpeL (21.8%) by efficiently providing pertinent guidelines for each state and minimizing the burden on context capacity. Refer to Figure 15 for a list of example states and guidelines.
Q3. How does AutoGuide perform when combined with test-time self-feedback approaches like Reflexion?
Our state-aware guidelines effectively provide inter-task knowledge by considering multiple tasks in offline data. Meanwhile, self-feedback methods (e.g., Reflexion) offer intra-task knowledge based on environmental feedback during test time. In this question, we explore the effectiveness of integrating both inter-task and intra-task information. The results presented in Table 3 demonstrate that the combination of AutoGuide with Reflexion achieves the highest performance in the WebShop benchmark. Hence, we find that our state-aware guidelines positively complement the intra-task knowledge of Reflexion. Another observation from Table 3 is that, while ExpeL + Reflexion outperforms ExpeL alone, this combination is not as effective as other approaches. This limitation may stem from ExpeL introducing irrelevant knowledge, potentially leading to conflicts with Reflexion’s feedback and having an adverse impact on the decision-making process.
Figure 5: Extracting state-aware guidelines when offline experiences have either success or failure trajectories. With minor modifications, AutoGuide can extract useful knowledge from offline data and outperform the ReAct baseline.
4.3 Analysis of AutoGuide
Q4. Can AutoGuide extract state-aware guidelines when offline data has either success or failure trajectories?
We explore how AutoGuide can be adapted to address extreme types of offline data, comprising either successful or failed trajectories. Intuitively, successful trajectories exhibit shared patterns of successful behaviors, whereas unsuccessful trajectories may fail for a variety of reasons. Based on this insight, when offline experiences contain exclusively successful instances, we sample multiple trajectories and prompt GPT-4-turbo to identify states exhibiting shared common behaviors across them. Then, for each state, AutoGuide applies the state summarization and guideline extraction modules to construct state-aware guidelines. In the case of failures only, we leverage failed trajectories and a one-shot example from ReAct as a successful demonstration to generate state-aware guidelines based on our modules.
The results are shown in Figure 5. Interestingly, even without having pairs of success and failure trajectories for the same task, the different variants of AutoGuide can extract valuable knowledge from offline data, leading to improvements in test-time performance compared to ReAct, which does not leverage any offline data. As such, our method is flexible and robust to the type of available offline data. However, in general cases, we expect to have both successful and failed trajectories, which provide the most direct insights for AutoGuide to extract useful guidelines.
Q5. How does each component of AutoGuide contribute to the final results?
We evaluate the impact of different components within AutoGuide on its performance in WebShop, as detailed in Table 4. We examine two variants: ReAct+SS and ReAct+GES. The ReAct+SS, which incorporates state summaries into observations without guidelines, shows improvement over ReAct. This suggests that state summaries enhance decision-making by verifying the current state before action selection. ReAct+GES, which generates guidelines from trajectories without state summaries and employs GPT-3.5-turbo for guideline selection, also enhances performance but is less effective than the full AutoGuide. This indicates that choosing relevant guideline based on the trajectory alone is more challenging than using state summary. Therefore, integrating both state summaries and guidelines is crucial for maximizing the benefits of AutoGuide.
[…]
Algorithm SS GES WebShop SR
@@ 4.3 Analysis of AutoGuide
Table 4: Ablation study of AutoGuide, analyzing each module’s contribution in the WebShop benchmark. SS denotes our state summarization module, and GES denotes the guideline extraction and selection modules.

### v2

@@ 3.3 Applying Context-Aware Guidelines at Test Time
After extracting a set of context-aware guidelines $\mathcal{G}$ from offline experiences, our method employs these guidelines to enhance the decision-making of an LLM agent during testing. At each timestep, AutoGuide identifies the context of the current test trajectory $\bm{\tau}$ up to timestep $t$ (which represents the agent’s interactions up to the current time-step) using our context identification module $\mathcal{M}_{\mathrm{context}}$. Our guideline selection module $\mathcal{M}_{\mathrm{select}}$ then selects relevant guidelines for context from $\mathcal{G}$. More specifically, the module applies context as the key to fetch a set of possible guidelines $\mathcal{G}[\textsc{context}]$. If there are more than $k$ guidelines in $\mathcal{G}[\textsc{context}]$, $\mathcal{M}_{\mathrm{select}}$ prompts an LLM to choose top-$k$ guidelines for the specific $\bm{\tau}$:
relevant $\begin{split}\textsc{relevant guidelines}\leftarrow\mathcal{M}_{\mathrm{select}}(\textsc{context},\bm{\tau};\mathcal{G},k),\end{split}$ (3)
@@
where Section C.4 details the prompt template for this selection procedure. Subsequently, AutoGuide incorporates both the context and relevant guidelines into the agent’s action generation prompt. Therefore, the agent selects an action by considering the provided context and guidelines (see Figure 1 for an example). This process iterates until the end of the test trajectory (see Algorithm 2).
@@
Key benefits of AutoGuide. First, the extraction of context-aware guidelines in AutoGuide offers the inherent benefit of providing relevant guidelines for the context of interest. This capability is important since neglecting the specific context in which a guideline applies can confuse the agent’s decision-making process. The second key benefit is the generation of concise natural language guidelines, which can be seamlessly incorporated into any prompt-based LLM agent. Lastly, AutoGuide generates guidelines at the individual context level rather than at the trajectory level. Given that a single incorrect action can lead to a complete failure, it is essential to provide detailed assistance in each action selection process. With these advantages, we demonstrate in the next section that our approach significantly enhances the performance of LLM agents.
Figure 3: Sequential decision-making benchmark domains considered in our work: ALFWorld [24], WebShop [33], WebArena [8], and multi-modal real-world websites. Graphic credit: [34, 33, 8, 35].
@@ 4 Evaluation
This section demonstrates the efficacy of AutoGuide by conducting experiments on a diverse suite of sequential decision-making benchmark domains. We also perform important analyses about AutoGuide, such as the ablation study of different AutoGuide components, comparison to in-context learning, and generalization to out-of-domain tasks. We refer to Appendix B for additional experimental details.
@@ 4.1 Evaluation Setup
4.1.1 Sequential Decision-Making Benchmark Domains
@@
We consider the following interactive sequential decision-making benchmarks to study various aspects of AutoGuide (see Figure 3):
@@
  • ALFWorld [24]: In this embodied benchmark, an LLM agent interacts with an environment to carry out household tasks, such as placing a pan on the dining table. Observations and actions are expressed in natural language statements, and the agent must navigate through the space and manipulate objects to successfully complete the tasks.
@@
  • WebShop [33]: This interactive web environment simulates the task of online shopping on an e-commerce website. The agent’s goal is to understand a text instruction and buy a product that meets specified criteria. This involves querying the website’s search engine, understanding the descriptions and details of each item, and selecting necessary options.
@@
  • WebArena [8]: This web-based benchmark introduces realistic environments by replicating the functionality and data found in popular web domains (e.g., Gitlab, Reddit, Wikipedia). Compared to WebShop, WebArena presents more challenges and difficulties for an LLM agent due to its large observation and action space, along with tasks that involve longer planning horizons. We focus oe the Reddit domain for the main WebArena experiments.
  • Real-world multi-modal websites: Finally, we consider evaluating AutoGuide on a variety of real-world website tasks. These span from a collaborative software development platform (e.g., GitHub) to a flight search engine (e.g., Google Flights) and an online education platform (e.g., Coursera). Please refer to Section B.4 for example tasks. In particular, in comparison to WebShop and WebArena, we design our tasks to be multi-modal such that the agent must consider both visual (e.g., images) and textual information (e.g., HTML) to complete these tasks.
Figure 4: Trajectories of ReAct, ExpeL, and AutoGuide from the same test task. ReAct (Left) chose the wrong item, consequently failing the task in the end. ExpeL (middle) was confused by guidelines that were irrelevant to current context, leading to incorrect reasoning and actions. AutoGuide (right) selects relevant guidelines to the agent’s context, enabling the agent to accomplish the task.
4.1.2 Baselines
We compare AutoGuide against the following baseline approaches to study the effect of context-aware guidelines (refer to Appendix B for more details):
  • ReAct [23]: This LLM-based planning method integrates reasoning and acting to address sequential decision-making tasks. However, it does not leverage offline experiences and thus suffers from the limited understanding of pre-trained LLMs in downstream domains.
  • ExpeL [31]: This method also extracts natural language knowledge from offline data. However, it fails to consider the applicability of guidelines and does not generate context-aware guidelines. Instead, it provides all guidelines to an LLM agent without filtering out irrelevant ones based on the current context. ExpeL has two contributions, the guideline generation and in-context example selection module. Because the latter is orthogonal to our analysis and can be seamlessly combined with our method, we consider ExpeL with guidelines in our experiments.
  • Reflexion [27]: This approach converts environmental feedback into text statements to assist an LLM agent (e.g., ReAct) in the next trial. The baseline generates valuable feedback about solving a specific test task. We demonstrate how context-aware guidelines derived by AutoGuide can be combined with the feedback.
4.1.3 Implementation
We collect offline experiences either by running ReAct and Reflexion, or incorporating human demonstrations. We use ReAct with GPT-3.5-turbo as our base LLM agent for WebShop and ALFWorld and GPT-4-turbo for WebArena. For each benchmark, we apply the same GPT model for action generation, context identification, and guideline selection. We extract context-aware guidelines from offline data with GPT-4-turbo and evaluate their effectiveness by applying them to the test set with non-overlapping tasks. We refer to Appendix B for more details.
[…]
Algorithm Offline data? Context aware? ALFWorld [24] WebShop [33] WebArena [8]
[…]
ReAct [23] 54.5% 66.4 30% 8.0%
ExpeL [31] 59.0% 60.9 35% 21.8%
AutoGuide 79.1% 73.4 46% 47.1%
ReAct [23] + Reflexion [27] 67.2% 77.1 51% N/A
ExpeL [31] + Reflexion [27] 71.6% 71.7 42% N/A
AutoGuide + Reflexion [27] 88.1% 81.4 57% N/A
@@ 4.1.3 Implementation
Table 1: Test reward and success rate on ALFWorld, WebShop, and WebArena. The base agent model for ALFWorld and WebShop is GPT-3.5-turbo and for WebArena is GPT-4-turbo. Reflexion is done by GPT-4-turbo for at most 3 trials. In our experiments, due to token limit of GPT, we did not experiment with Reflexion on WebArena tasks.
4.2 Main Results
Q1. How effective is AutoGuide compared to baselines without context-aware guidelines?
To answer this question, we compare methods on ALFWorld, WebShop, and WebArena benchmarks. The performance on the test datasets is presented in Table 1. There are three notable observations:
  1. 1.
    Effectiveness of context-aware guidelines. Our approach surpasses baseline performance in both ALFWorld and WebShop, achieving the highest test rewards and success rates in Table 1. These results highlight the effectiveness of employing context-aware guidelines in language-based decision-making domains. To further examine the action selection differences among ReAct, ExpeL, and our method, we present their trajectories in Figure 4. We observed that ReAct makes common mistakes such as trying to take soapbar that is not visible, or taking a soapbottle instead of soapbar due to their similar names. Both ExpeL and AutoGuide improve on this by extracting guidelines from similar mistakes in the offline experience. However, ExpeL often erroneously applies incorrect guidelines due to the availability of all guidelines at each timestep. In Figure 4, ExpeL mistakenly attends to the second guideline “ensure to specify the item’s number and location …”, leading to wrong reasoning and action. AutoGuide presents relevant guidelines at necessary moments, enabling accurate task completion by avoiding the mistakes seen in ExpeL and ReAct.
  2. 2.
    Importance of providing pertinent knowledge. ExpeL approach helps ReAct by extracting knowledge from offline experiences, but its impact is not as significant as AutoGuide. Recall that for ExpeL, the guidelines are neither generated for specific contexts at training time nor selected to only provide context-aware guidelines at test time. As a result, irrelevant guidelines can be introduced to an agent, potentially causing confusion for the agent. Consequently, the result highlights the significance of providing relevant guidelines conditioned on contexts for LLM agents.
  3. 3.
    Scalability to complex environments. We conduct experiments on WebArena-Reddit, which features more diverse tasks on realistic and complex websites requiring longer action sequences. This domain has a larger observation space and a more complex action space (e.g., scrolling). Table 1 presents the results, where AutoGuide achieves the highest success rate with a significant margin when compared to ReAct and ExpeL. We observe that ReAct scores low task success rate (8.0%) in WebArena due to the complex observation and action spaces and longer task horizon. In ExpeL, the issue of presenting all guidelines to an agent is exacerbated in the WebArena compared to simpler environments like ALFWorld and WebShop. WebArena’s wide variety of tasks across different domains requires a larger number of guidelines to cover the knowledge needed for all tasks and domains. This results in either an overload of irrelevant guidelines that could mislead the agent or a lack of crucial information when the number of guidelines is limited, as suggested in ExpeL [31]. In contrast, AutoGuide achieves a more significant performance enhancement (47.1%) compared to ExpeL (21.8%) by efficiently providing pertinent guidelines and minimizing the burden on context capacity. We refer to Figure 14 for a list of example contexts and guidelines.
[…]
Algorithm GitHub Flights Coursera
[…]
SoM [36] 2/30 5/20 1/20
AutoGuide 19/30 9/20 14/20
Table 2: Test results of AutoGuide on 3 real-world web domains within multi-modal settings. The base agent model runs with GPT-4V and applying context-aware guidelines significantly improves the performance.
[…]
Algorithm WebShop
[…]
ReAct (1-shot) 66.4 30%
ReAct (2-shot) 66.0 35%
ReAct (4-shot) 70.2 37%
ReAct (6-shot) 71.0 38%
AutoGuide 73.4 46%
Table 3: Analysis of AutoGuide against ReAct with varying numbers of in-context examples.
[…]
Top-$k$ WebShop
[…]
$k\!=\!0$ 30%
$k\!=\!1$ 42%
$k\!=\!2$ 46%
$k\!=\!3$ 47%
$k\!=\!5$ 43%
Table 4: Ablation study of AutoGuide using various top-$k$ values.
Q2. How does AutoGuide perform when combined with test-time self-feedback approaches?
Our context-aware guidelines effectively provide inter-task knowledge by considering multiple tasks in offline data. Meanwhile, self-feedback methods (e.g., Reflexion) offer intra-task knowledge based on environmental feedback during test time. In this question, we explore the effectiveness of integrating both inter-task and intra-task information. The results presented in Table 1 demonstrate that the combination of AutoGuide with Reflexion achieves the highest performance in the WebShop and ALFWorld benchmarks. Hence, we find that our context-aware guidelines positively complement the intra-task knowledge of Reflexion. Another observation from Table 1 is that, while ExpeL + Reflexion outperforms ExpeL alone, this combination is not as effective as other approaches. This limitation may stem from ExpeL introducing irrelevant knowledge, potentially leading to conflicts with Reflexion’s feedback and having an adverse impact on the decision-making process.
Q3. Can AutoGuide generate context-aware guidelines for multi-modal inputs?
Going beyond text-only inputs is an essential step toward building capable agents for solving real-world environments and tasks. We test AutoGuide in a complex multi-modal setting, where each observation includes image and text information. Specifically, we introduce a set of real-world website navigation tasks in 3 domains: GitHub, Google Flights, and Coursera. For these multi-modal tasks, we employ the Set-of-Marks (SoM) agent [36, 5] as our base method. The SoM prompting improves the visual grounding capabilities of large multi-modal models such as GPT-4V by adding visually distinguishable marks to image inputs [36]. We apply AutoGuide with GPT-4V to generate natural language context-aware guidelines from collected trajectories with both image and text observations. Table 4 shows the effectiveness of AutoGuide, demonstrating its generalization ability to complex real-world multi-modal settings. We refer to Figure 15 for example context-aware guidelines.
4.3 Analyses of AutoGuide
Q4. How does AutoGuide compare to ReAct with varying numbers of in-context examples?
Table 4 shows that, while increasing the number of in-context examples for ReAct gradually improves performance, there is a plateau at a certain number of shots. Additionally, ReAct with more than 6 shots often exceeds the token limit of GPT-3.5-turbo. These results indicate that directly inputting raw trajectories into ReAct for in-context learning is not an effective way to fully leverage offline data. In contrast, AutoGuide extracts knowledge from entire training trajectories by summarizing them into concise context-aware guidelines, making them easy to integrate with prompt-based agents.
Q5. How does altering the number of top-$k$ guidelines impact the performance of AutoGuide?
We conducted an ablation study on WebShop using various values of $k$ in Table 4. We find that employing context-aware guidelines consistently outperforms the no-guideline baseline ($k\!=\!0$; ReAct). The $k=3$ yields the best performance. The largest $k$ value of 5 can lead an LLM agent to overthink, potentially resulting in a slight decrease in performance. Conversely, a smaller $k$, like $k\!=\!1$, may cause LLM to overlook additional helpful guidelines, leading to slightly worse performance.
Q6. How do AutoGuide’s context-aware guidelines generalize to out-of-domain environments?
We conduct an experiment to further demonstrate AutoGuide’s out-of-domain capability across different domains but relevant tasks. We extract context-aware guidelines from WebShop and apply them to WebArena-Shopping, which is a distinct domain with variations in observation/action spaces, task intentions, and episodic horizons. For this domain adaptation case, we additionally incorporate a grounding module to align the context-aware guidelines from WebShop to WebArena’s observations based on GPT-4-Turbo. As shown in Table 6, the transferred guidelines bring a notable improvement in success rates compared to the ReAct baseline in WebArena Shopping.
Q7. How does each component of AutoGuide contribute to the final results?
We evaluate the impact of different components within AutoGuide on its performance in WebShop, as detailed in Table 6. We examine two variants: ReAct+CI and ReAct+GES. The ReAct+CI, which incorporates contexts into observations without guidelines, shows improvement over ReAct. This suggests that contexts enhance decision-making by verifying the current state before action selection. ReAct+GES, which generates guidelines from trajectories without contexts and employs GPT-3.5-turbo for guideline selection, also enhances performance but is less effective than the full AutoGuide. This indicates that choosing relevant guidelines based on the trajectory alone is more challenging than using contexts. Therefore, integrating both context summaries and guidelines is crucial for maximizing the benefits of AutoGuide.
[…]
Algorithm WebArena–Shopping
[…]
ReAct 10.2%
AutoGuide 20.4%
Table 5: Out-of-distribution generalization of context-aware guidelines from WebShop on the 98 WebArena–Shopping tasks that have a product in the intent template.
@@ 4.3 Analyses of AutoGuide
Table 6: Ablation study of AutoGuide, analyzing each module’s contribution in WebShop. CI denotes our context identification module, and GES denotes the guideline extraction and selection modules.